# Random coverage from within with variable radii, and Johnson-Mehl cover timesthanks: Supported by EPSRC grant EP/T028653/1

### v2

@@ 2.2 Results for the Johnson-Mehl model
A series of videos of very high intensity restricted Johnson-Mehl processes inside $A=[0,1]^{2}$ can be viewed here [^3] https://www.youtube.com/playlist?list=PLiaV5rk6Gk7rh6nvAts1w4WZrHE0TKeW2 An interactive tool for generating Johnson-Mehl tessellations is available at https://frankiehiggs.pyscriptapps.com/johnson-mehl-plot/latest. The cells are coloured (with five different colours) so that no two adjacent cells in the final “tessellation” have the same colour.
To illustrate the above remark about which point is covered last when $d=2$, we have included one video in the playlist where the last point to be covered is at the boundary, and one where it is in the interior.
Results of more simulations, comparing the distribution of cover times for the restricted and unrestricted J-M models with finite $\rho$ and their limiting distributions, are presented in Figure 1.
@@
Suppose we take $d=2$ and $A=[0,1]^{2}$. Comparing (2.11) with (2.13) we see that the limiting distribution of $T_{\rho}^{3}$ (appropriately scaled and centred) is different from that of $\tilde{T}_{\rho}^{3}$ (scaled and centred the same way). Similarly, comparing (2.12) with (2.14) we see that the limiting distribution of $\tau_{L}^{3}$ (appropriately scaled and centred) is different from that of $\tilde{\tau}_{L}^{3}$ (scaled and centred the same way).
When $d\geq 3$ and $\partial A\in C^{2}$, we see from (2.11) and (2.16) that even the centring constants required to get a nondegenerate limit distribution for $\omega_{d}\rho T_{\rho}^{d+1}$ are different (larger) compared to the corresponding constants for $\omega_{d}\rho\tilde{T}_{\rho}^{d+1}$. Again similar remarks apply for $\tau_{L}$ and $\tilde{\tau}_{L}$.
@@
As mentioned earlier, [1] considers cover times for more general J-M models where the growth rate of cells, or the arrival rates of seeds, are not necessarily constants. If growth rates are non-constant, then it is rather problematic to model the covered region at a given time as simply a SPBM; one has to modify our initial description of the cell growth model to allow a faster-growing cell to pass through a slower-growing cell and come out the other side. If one is willing to do this (which seems to be the approach in [1]), the methods of this paper should be applicable for dealing with boundary effects for these more general J-M models too.
@@
The distribution of the maximum of two independent Gumbel variables with different scale parameters is known as a two-component extreme value (TCEV) distribution in the hydrology literature [12].
@@
Theorems 2.9 and 2.10 tell us that when $d=2$ and $A$ is polygonal or $\partial A\in C^{2}$, the random variables $\omega_{d}\rho T_{\rho}^{d+1}$ and of $\omega_{d}\tau_{L}^{d+1}$, suitably centred, are asymptotically TCEV distributed, while if $d\geq 3$ and $\partial A\in C^{2}$, they are asymptotically Gumbel distributed.
@@ 2.3 Strategy of proof
Next, we briefly describe the strategy for the proof, in Sections 3 and 4, of the weak convergence results that were stated above.
We shall use a known result (Lemma 3.2) giving the limiting probability of covering a bounded region of $\mathbb{R}^{d}$ by an SPBM on the whole of $\mathbb{R}^{d}$, in the limit of high intensity $n$ and small balls of random radius, i.e. distributed as $r_{n}Y$ for a specified nonnegative random variable $Y$ and for constants $r_{n}$ that become small as $n$ becomes large.
We shall consider for large $n$ and suitable choice of $r_{n}$, the SPBM with balls of radii distributed as $r_{n}Y$, restricted to a $d$-dimensional half-space $\mathbb{H}$. In Lemma 3.5 we determine the limiting probability that a given bounded set within the hyperplane $\partial\mathbb{H}$ is covered, by applying Lemma 3.2 in $d-1$ dimensions. Moreover we will show that the probability that a region in the half-space within distance $n^{\zeta}r_{n}$ of that set is covered with the same limiting probability, where ${\zeta}$ is a small positive constant.
@@
In the case where $Y$ is unbounded, to obtain the required independence, for example between coverage events for different faces of our polyhedron or polygon, we consider only those balls of radius at most $n^{\zeta}r_{n}$, with a separate argument (Lemmma 3.3) to show this does not affect the limiting coverage probabilities.
Applying Lemma 3.2 for an SPBM with $Y$ uniformly distributed on $[0,1]$ yields a proof of Proposition 2.8. Similarly, applying Theorems 2.1 and 2.2 for an SPBM with $Y$ uniformly distributed on $[0,1]$ yields proofs of Theorems 2.9 and 2.10 respectively.
3 Proof of Theorems 2.1 and 2.2
3.1 Preliminaries
Given two sets $\mathcal{X},{\cal Y}\subset\mathbb{R}^{d}$, we set $\mathcal{X}\triangle{\cal Y}:=(\mathcal{X}\setminus{\cal Y})\cup({\cal Y}\setminus\mathcal{X})$, the symmetric difference between $\mathcal{X}$ and ${\cal Y}$. Also, we write $\mathcal{X}\oplus{\cal Y}$ for the set $\{x+y:x\in\mathcal{X},y\in{\cal Y}\}$, and #$\#(\mathcal{X})$ for the number of elements of $\mathcal{X}$ (possibly +$\infty$+ ∞). Given also $x\in\mathbb{R}^{d}$ we write $x+{\cal Y}$ for $\{x\}\oplus{\cal Y}$.
From time to time we shall use the following geometrical lemma.
Lemma 3.1 (Geometrical lemma).
$|B(x,r)\cap A|\geq((\omega_{d}/2)-\varepsilon)r^{d},~{}~{}~{}~{}\forall x\in A,r\in(0,r_{0}).$ (3.1)
Proof.
See [4, Lemma 3.2(a)]
We shall repeatedly use the following result from [10], which is based on results in Janson [5] or Hall [3]. Recall that $c_{d}$ and $Z_{A}(n,Y)$ were defined at (2.2) and just before Theorem 2.1 respectively.
@@ 3.1 Preliminaries
$\lim_{\lambda\to\infty}\left(\alpha\delta(\lambda)^{d}\lambda-\log\lambda-(d+k-2)\log\log\lambda\right)=\beta.$ (3.2)
@@
$\lim_{\lambda\to\infty}\mathbb{P}[B_{\lambda}\subset Z_{\mathbb{R}^{d},k}(\lambda,\delta(\lambda)Y)=\exp\left(-\left(\frac{c_{d}(\mathbb{E}\,[Y^{d-1}])^{d}}{(k-1)!(\mathbb{E}\,[Y^{d}])^{d-1}}\right)|B|e^{-\beta}\right).$ (3.3)

### v3

@@ 2.1 Results for the Johnson-Mehl model
A series of videos of very high intensity restricted Johnson-Mehl processes inside $A=[0,1]^{2}$ can be viewed here [^2] https://www.youtube.com/playlist?list=PLiaV5rk6Gk7rh6nvAts1w4WZrHE0TKeW2. The cells are coloured (with five different colours) so that no two adjacent cells in the final “tessellation” have the same colour. To illustrate the above remark about which point is covered last when $d=2$, we have included one video in the playlist where the last point to be covered is at the boundary, and one where it is in the interior.
An interactive tool for generating Johnson-Mehl tessellations is available at https://frankiehiggs.pyscriptapps.com/johnson-mehl-plot/latest.
The results of more simulations, comparing the distribution of cover times for the restricted and unrestricted J-M models with finite $\rho$ and their limiting distributions, are presented in Figure 2.
In [3], it is suggested that the same limit theorem should apply to $\tau_{L}$ as to $\tilde{\tau}_{L}$, as the boundary effects would not affect the limit. This would imply that the same limit should apply to $T_{\rho}$ as to $\tilde{T}_{\rho}$. Our Theorem 2.2 shows that in fact the boundary effects do affect the limiting distribution of $g(T_{\rho},\rho)$, and the simulations shown in Figure 2 back this up.
@@
Suppose we take $d=2$ and $A=[0,1]^{2}$. Comparing (6) with (8) we see that the limiting distribution of $T_{\rho}^{3}$ (appropriately scaled and centred) is different from that of $\tilde{T}_{\rho}^{3}$ (scaled and centred the same way). Similarly, comparing (7) with (9) we see that the limiting distribution of $\tau_{L}^{3}$ (appropriately scaled and centred) is different from that of $\tilde{\tau}_{L}^{3}$ (scaled and centred the same way).
When $d\geq 3$ and $\partial A\in C^{2}$, we see from (6) and (11) that even the centring constants required to get a nondegenerate limit distribution for $\omega_{d}\rho T_{\rho}^{d+1}$ are different (larger) compared to the corresponding constants for $\omega_{d}\rho\tilde{T}_{\rho}^{d+1}$. Again similar remarks apply for $\tau_{L}$ and $\tilde{\tau}_{L}$.
@@
As mentioned earlier, in [3] Chiu considers cover times for more general J-M models where the growth rate of cells, or the arrival rates of seeds, are not necessarily constants. If growth rates are non-constant, then it is rather problematic to model the covered region at a given time as simply a SPBM; one has to modify our initial description of the cell growth model to allow a faster-growing cell to pass through a slower-growing cell and come out the other side. If one is willing to do this (which seems to be the approach in [3]), the methods of this paper should be applicable for dealing with boundary effects for these more general J-M models too.
These more general J-M type models with non-constant growth rates (in particular parabolic growth rates) and non-constant arrival rates, and with the balls allowed to interpenetrate, have been studied in the literature in relation to the understanding of asymptotics of convex hulls of random points [19], maximal points of a multidimensional sample [19] and constructing generalized Delaunay tessellations [5]. It could be interesting to try to map our methods here onto questions of interest in these related models; this is a possible future direction of research.
@@
The distribution of the maximum of two independent Gumbel variables with different scale parameters (e.g. the variable shown at (13)) is known as a two-component extreme value (TCEV) distribution in the hydrology literature [18].
@@
Theorems 2.2 and 2.3 tell us that when $d=2$ and $A$ is polygonal or $\partial A\in C^{2}$, the random variables $\omega_{d}\rho T_{\rho}^{d+1}$ and of $\omega_{d}\tau_{L}^{d+1}$, suitably centred, are asymptotically TCEV distributed, while if $d\geq 3$ and $\partial A\in C^{2}$, they are asymptotically Gumbel distributed.
The TCEV arises elsewhere in geometric extreme value theory. For example, suppose $L_{n,k,d}$ denotes the largest $k$-nearest neighbour link in a sample of $n$ uniform random points in $0,1]^{d}$, $n>k$. That is, denoting these points $p_{1},\ldots,p_{n}$, we have $L_{n,k,d}=\max_{1\leq i\leq n}(k$-$\min_{j\leq d,j\neq n}\|p_{i}-p_{j}\|)$, where $k$-$\min(\cdot)$ means the $k$th smallest of a set of at least $k$ numbers. It is shown in [16] that $n\pi L_{n,2,2}^{2}$, suitably centred, is asymptotically TCEV distributed. Again this is due to boundary effects, and again, these have sometimes been missed in the literature; for example it is claimed in [12, p214], incorrectly, that $n\pi L_{n,2,2}^{2}$ suitably centred is asymptotically Gumbel. The limiting distribution of $L_{n,k,d}$ (suitably transformed) for general fixed $(k,d)$ is given in [13, Theorem 8.4], which the authors of [12] were apparently unaware of.
2.2 Results for the spherical Poisson Boolean model
As discussed in the introduction, we shall derive our results for the J-M model from results on coverage by the restricted SPBM, which are of independent interest and which we now state. Given a nonnegative random variable $Y$ and given $n>0$, suppose we have a collection of balls of independent random radii with the distribution of $Y$, centred on the points of a homogeneous Poisson point process in $A$ of intensity $n$.
Theorem 2.9 (Limiting probability of $k$-coverage of a polygonal domain by a restricted SPBM).
$\mathbb{P}[A\subset Z_{A,k}(n,r_{n}Y)]\to\exp\Big(-\Big(\frac{(\mathbb{E}\,[Y])^{2}}{\mathbb{E}\,[Y^{2}]}\Big){\bf 1}_{\{k=1\}}|A|e^{-\beta}-\Big(\frac{c_{2,k}\mathbb{E}\,[Y]|\partial A|}{(\mathbb{E}\,[Y^{2}])^{1/2}}\Big)e^{-\beta/2}\Big),$ (14)
where we set $c_{2,k}:=\pi^{-1/2}(1/2)^{k-1}/(k-1)!$.
For general $d,k\in\mathbb{N}$ with $d\geq 2$ we define $c_{d,k}$ (as in [15]) by
$c_{d,k}$ $:=c_{d-1}\omega_{d}^{2-d-1/d}\omega_{d-1}^{2d-3}\omega_{d-2}^{1-d}(1-1/d)^{d+k-3+1/d}2^{-1+1/d}/(k-1)!.$ (15)
It can be checked that
$c_{d,1}=((d-1)!)^{-1}2^{1-d}(d-1)^{d-2+1/d}\pi^{(d/2)-1}\Gamma\left(\frac{d+1}{2}\right)^{1-d}\Gamma\left(\frac{d}{2}\right)^{d-1+1/d},$ (16)
and $c_{d,k}=c_{d,1}(1-1/d)^{k-1}/(k-1)!$. Also $c_{2,1}=\pi^{-1/2}$ and $c_{3,1}=2^{-4}\pi^{5/3}$. In the right hand side of (16) the first factor of $((d-1)!)^{-1}$ was given, incorrectly, as $(d!)^{-1}$ in [15].
Theorem 2.10 (Limiting probability of $k$-coverage of a smoothly bounded region in $d\geq 2$ by a restricted SPBM).
$n\omega_{d}r_{n}^{d}\mathbb{E}\,[Y^{d}]-(2-2/d)\log n-2(d+k-3+1/d)\log\log n\to\beta$ (17)
as $n\to\infty$. If $d=2$ then (14) holds, while if $d\geq 3$ then
$\lim_{n\to\infty}\mathbb{P}[A\subset Z_{A,k}(n,r_{n}Y)]=\exp\Big(-c_{d,k}\Big(\frac{(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(\mathbb{E}\,[Y^{d}])^{d-2+1/d}}\Big)|\partial A|e^{-\beta/2}\Big).$ (18)
Remark 2.11.
Taking $Y\equiv 1$ in these results gives Theorem 3.2 and Theorem 3.1 of [15]. That paper also includes parallel results (in the case $Y\equiv 1$) for a situation where the number of grains in the spherical Boolean model is a large deterministic constant rather than a Poisson variable. We expect that a result along these lines could be given for general $Y$ using the methods of this paper.
Remark 2.12.
The condition $\overline{A^{o}}=A$ should have been included in the statement of [15, Theorem 3.1]. For example, if $d=2$ and $A$ is the union of a disk (filled in) and a circle (not filled in, and disjoint from the disk) then that theorem does not apply.
Remark 2.13.
Remark 2.14.
If $\mathbb{E}\,[Y^{d}]=\infty$ then $\mathbb{P}[A\subset Z_{\mathbb{R}^{d},k}(n,rY)]=1$ for any $k\in\mathbb{N}$ and any $r>0$; see e.g. [9, Theorem 16.4] for the case $k=1$. However, one could still look for a sequence $(r_{n})$ such that $\mathbb{P}[A\subset Z_{A,1}(n,r_{n}Y)]$ converges to a nontrivial limit. It seems plausible in this case that $A$ is most likely to be covered (if at all) by a small number of balls. This looks like an interesting problem that lies beyond the scope of this paper.
Remark 2.15.
A series of videos illustrating the SPBM and Remark 2.14 can be viewed at this link [^3] https://www.youtube.com/playlist?list=PLiaV5rk6Gk7qEXpLOU7FSvN4b8dy1_GJn. There are three videos, with $n=1000$, $n=10000$ and $n=20000$ respectively. Each video has a Poisson point process of intensity $n$ inside $0,1]^{2}$ with points $\{p_{1},\dots,p_{N_{n}}\}$. The frame at time $t$ is an image of a tessellation corresponding to a restricted SPBM, $Z_{A,1}(n,r_{t}Y_{t})$, where $A=[0,1]^{2}$, $Y_{t}$ is Pareto distributed with shape parameter $\alpha(t)$ and $r_{t}$ is chosen so that $A\subset Z_{A,1}(n,r_{t}Y_{t})$. There are $N_{n}$ cells in the frame: the pixel with centre $x\in[0,1]^{2}$ is assigned to a cell based on $\mathrm{argmin}_{i}\inf\{r\geq 0:x\in B(p_{i},rY_{t,i})\}$, where $(Y_{t,i})_{1\leq i\leq N_{n}}$ are independent copies of $Y_{t}$. In other words, for $t$ fixed the tessellation is generated by a growth process for which all seeds are born at time 0, growth rates are random and distributed like $Y_{t}$, and the growing cells are interpenetrable.
The videos after the point where $\mathbb{E}\,[Y_{t}^{2}]=\infty$ illustrate Remark 2.14. In the period when $\mathbb{E}\,[Y_{t}^{2}]=\infty$ but $\mathbb{E}\,[Y_{t}^{7/4}]<\infty$, there is not yet a single cell containing more than half the area of $0,1]^{2}$, indicating that there may be a non-trivial period where several balls are required to cover $0,1]^{2}$.
@@ 2.3 Strategy of proof
Next, we briefly describe the strategy for the proof, in Sections 3 and 4, of the weak convergence results that were stated above. In both sections, we shall use a known result (Lemma 3.1) giving the limiting probability of covering a bounded region of $\mathbb{R}^{d}$ by an unrestricted SPBM, in the limit of high intensity $n$ and small balls of random radius, i.e. distributed as $r_{n}Y$ for a specified nonnegative random variable $Y$ and for constants $r_{n}$ that become small as $n$ becomes large.
We claim that the question of coverage for the (restricted) SPBM with uniformly distributed radii maps onto the same question for the restricted Johnson-Mehl model. Indeed, given $t>0$, consider a SPBM in $A$ with Poisson intensity $n$ and radii uniformly distributed over $[0,t]$. Label the centres $x_{1},\ldots,x_{N}$ with associated radii $t_{1},\ldots,t_{N}$. Then for each $i$, $t_{i}$ is uniformly distributed over $[0,t]$ and therefore so is $t-t_{i}$. Hence by the Marking theorem (see e.g. [9]), the point process $\eta:=\{(x_{i},t-t_{i}),1\leq i\leq N\}$ is a Poisson process in $A\times[0,t]$ with intensity given by the product of Lebesgue measure on $A$ and the uniform probability distribution on $[0,t]$, i.e. $n/t$ times Lebesgue measure on $A\times[0,t]$. Hence the covered region for the original SPBM in $A$ is the same as the covered region for the restricted J-M model (run up to time $t$) obtained by using the Poisson process $\eta$. We can argue similarly in the unrestricted case too.
Using this claim and applying Lemma 3.1 for the SPBM with $Y$ uniformly distributed on $[0,1]$ will yield a proof of Proposition 2.1. Similarly, applying Theorems 2.9 and 2.10 with uniform $Y$ will yield proofs of Theorems 2.2 and 2.3 respectively.
Our strategy for proving Theorems 2.9 and 2.10 goes as follows. We shall consider, for large $n$ and suitable choice of $r_{n}$, the SPBM with radii distributed as $r_{n}Y$, restricted to a $d$-dimensional half-space $\mathbb{H}$. In Lemma 4.4 we determine the limiting probability that a given bounded set within the hyperplane $\partial\mathbb{H}$ is covered, by applying Lemma 3.1 in d1d-1 dimensions. Moreover we will show that the probability that a region in the half-space within distance $n^{\zeta}r_{n}$ of that set is covered with the same limiting probability, where $\zeta$} is a small positive constant.
@@
In the case where $Y$ is unbounded, to obtain the required independence, for example between coverage events for different faces of our polyhedron or polygon, we consider only those balls of radius at most $n^{\zeta}r_{n}$, with a separate argument (Lemma 4.2) to show this does not affect the limiting coverage probabilities.
3 Proof of results for the J-M model
We now use Theorems 2.9 and 2.10 to prove Theorems 2.2 and 2.3. The proof of Theorems 2.9 and 2.10 is much longer and we defer this to the next section.
We shall repeatedly use the following result from [15], which is based on results in [8] or [6]. Recall that $c_{d}$ and $Z_{A}(n,Y)$ were defined at (1) and just before Theorem 2.9 respectively.
@@ 3 Proof of results for the J-M model
$\lim_{\lambda\to\infty}\left(\alpha\delta(\lambda)^{d}\lambda-\log\lambda-(d+k-2)\log\log\lambda\right)=\beta.$ (19)
@@
$\lim_{\lambda\to\infty}\mathbb{P}[B_{\lambda}\subset Z_{\mathbb{R}^{d},k}(\lambda,\delta(\lambda)Y)=\exp\left(-\left(\frac{c_{d}(\mathbb{E}\,[Y^{d-1}])^{d}}{(k-1)!(\mathbb{E}\,[Y^{d}])^{d-1}}\right)|B|e^{-\beta}\right).$ (20)